# Dynamic link switching induces stable synchronized states in sparse networks

### v1

@@ III.1 Link-switching-induced inter-layer synchronization
We now demonstrate the role of switching the existing inter-layer links in establishing synchronization between the layers. Naturally, the time interval during which the links remain active before switching, referred to as the switching time $T_{swt}$, plays a central role in this process (problem with the concept ?). To explore this role, we first analyze the time evolution of the double-layer system for different values of $T_{swt}$ in a network with $N=400$ FHN oscillators per layer and only $N_{IL}=100$ inter-layer links connecting the mirror nodes between the layers.
The parameters used for the figures are all common and as following: $\varepsilon=0.05$, $a_{ij}=0.87$ (even $j$), $a_{ij}=0.97$ (odd $j$), $\phi=\frac{\pi}{2}-0.1$, $\sigma_{1}=\sigma_{2}=0.1$.
@@
In Fig. 2(a), for $T_{swt}=120$, the red curve represents the time evolution of the Euclidean distance between mirror nodes numbered $35$, defined as $E_{35}=\sqrt{(u_{i=1j=35}-u_{i=2j=35})^{2}+(v_{i=1j=35}-v_{i=2j=35})^{2}}$. In this figure, the time intervals shown in gray correspond to the presence of an inter-layer link between the mirror nodes $35$, while the time intervals shown in white indicate the absence of such a link. We observe that the distance $E_{35}$ exhibits high-amplitude oscillations during the time intervals marked in white, corresponding to the absence of an inter-layer link. In contrast, during the gray intervals, where an inter-layer link is present, $E_{35}$ oscillates with a significantly lower amplitude. Notice that, for $T_{swt}=120$, immediately after the gray time intervals, regardless of their duration, the amplitude of $E_{35}$ begins to increase as soon as the inter-layer link between mirror nodes $35$ is switched off. This behavior suggests that switching the inter-layer links at every $T_{swt}=120$ does not result in inter-layer synchronization. On the other hand, in Fig. 2(b), for $T_{swt}=23$, where the resulting time intervals marked in gray are significantly shorter, we observe a transient phase in which high and low amplitude oscillations alternate depending on the presence or absence of an inter-layer link. However, in this case, right after one gray time interval, the distance $E_{35}$ drops to zero, marking the end of the transient phase and indicating synchronization between the mirror nodes $35$. Finally, we emphasize that, although all links are switched periodically at every $T_{swt}$, a link connecting a specific pair of mirror nodes may be reassigned to the same pair after switching. Consequently, as observed in Fig. 2(a) and Fig. 2(b), the time intervals marked in gray can vary in size. Moreover, since $N_{IL}\ll N$, a given pair of mirror nodes may not receive an inter-layer link even after several switches, resulting in the time intervals marked in white also varying in size.
Now, we complement the analysis presented in Figs. 2(a) and 2(b) by verifying their corresponding behavior with respect to inter-layer synchronization. To achieve this, we first define a measure to estimate the level of inter-layer synchronization at each time instant as:
$E^{12}(t)=\frac{1}{N}\|\mathbf{x}_{1}(t)-\mathbf{x}_{2}(t)\|,$ (4)
@@
Figure 2 illustrates the dynamics of the double-layer system influenced by the switching of inter-layer links. We observe that switching the existing links at every $T_{swt}=23$ drives the system to inter-layer synchronization. To verify the robustness of this switching process, we evaluate the duration of the transient time required for the link switching to bring the layers into complete inter-layer synchronization.
In Fig. 3, we present the duration of the transient time required for the link switching to establish complete inter-layer synchronization, $T_{sync}$, as a function of the layer size $N$ for different switching times $T_{swt}=5,10,15,23,25,35,50$, while keeping the ratio $N_{IL}/N=0.25$ constant complying with the sparse network requirement. The transient time $T_{sync}$ increases monotonically with increasing $N$ for all switching times, demonstrating the effectiveness of the link switching procedure.
@@ III.2 Stability analysis of synchronized states
In the previous section, we observed that with a low density of inter-layer links, the dynamic redistribution of existing links among the FHN oscillators can promote the emergence of synchronous dynamics across the layers. More interestingly, the timing of these redistributions, defined by the switching time $T_{swt}$, plays a crucial role. However, whether the emergent synchronous dynamics is asymptotically stable remains an open question. To address this, we perform a transverse stability analysis using the master stability function (MSF) formalism on a simplified system consisting of two layers.
To implement the MSF in such a system, we treat each entire layer, comprising two oscillators, as a single higher-dimensional dynamical system, rather than as a network of lower-dimensional systems. Using this approach, Eq. (1) can be rewritten as:
@@
where $\mathbf{L}_{i}$ with $i=1,2$ represents the dynamical state of the entire layer, which comprises $N=2$ FHN oscillators labeled as $j=1,2$. The function $\mathbf{G}(\mathbf{L}_{i})$ encapsulates both the local dynamics of the FHN oscillators defined in Eq. (2) and the intra-layer coupling prescribed by the matrices $\mathcal{L}$ and $\mathcal{H}$, along with the parameter $\sigma_{i}$. The complete synchronized inter-layer state in the double-layer system occurs on a synchronization manifold $\mathbf{S}$ which is defined as $\mathbf{S}=\mathbf{L}_{1}=\mathbf{L}_{2}$. To assess the local stability of $\mathbf{S}$ using the MSF formalism, we introduce infinitesimal perturbations $\delta\mathbf{L}_{i}$ in all directions of the layer state space. This approach leads to the following variational equation governing the time evolution of $\delta\mathbf{L}_{i}$:
$\dot{\delta{\bf L\;}}_{i}=\left[\mathbf{DG}(\mathbf{S})+\sigma_{12}(\mathcal{L}^{I}\otimes\mathbf{D}\mathbf{\Gamma}(\mathbf{S},t))\right]\cdot\delta\mathbf{L}_{i},$ (6)
where $\mathbf{DG}(\mathbf{S})$ and $\mathbf{D\Gamma}(\mathbf{S})$ represent the Jacobian matrices of the intra-layer dynamics and the inter-layer coupling functions, respectively, evaluated at the synchronization manifold $\mathbf{S}$. The Laplacian matrix $\mathcal{L}^{I}$ is diagonalized with real eigenvalues $\lambda_{i}$ for $i=1,2$. Consequently, Eq. (6) becomes block diagonalizable, yielding a decoupled variational equation:
$\delta\dot{\mathbf{\Theta}}_{i}=\left[\mathbf{DG}(\mathbf{S})-\sigma_{12}|\lambda_{i}|\mathbf{D}\mathbf{\Gamma}(\mathbf{S},t)\right]\cdot\delta\mathbf{\Theta}_{i},$ (7)
@@
$\delta\dot{\mathbf{\Theta}}=\left[\mathbf{DG}(\mathbf{S})-2\sigma_{12}\mathbf{D}\mathbf{\Gamma}(\mathbf{S},t)\right]\cdot\delta\mathbf{\Theta}.$ (8)
The largest Lyapunov exponent of the system in Eq. (8) can be represented by the MSF $\Psi$. Since the coupling intensity $\sigma_{12}$ and the switching time $T_{swt}$ are the parameters of interest, we obtain the MSF as function of these parameters $\Psi(\sigma_{12},T_{swt})$.
Now, to demonstrate that a single inter-layer link is insufficient to establish synchronization between the oscillators of the two-layer system, we calculate $\Psi(\sigma_{12},T_{swt})$ as a function of $\sigma_{12}$ for the static case, i.e., without link switching. Hence, in Fig. 5, the blue curve illustrates the behavior of $\Psi(\sigma_{12})$ for the inter-layer link connecting oscillator number 1 from each layer (see schematic in Fig. 6(a) for the oscillators numbers). In this case, $\Psi(\sigma_{12})$ is positive within the considered interval of $\sigma_{12}$, indicating that the completely synchronized state is unstable for this inter-layer topology. Similarly, the red curve on the positive side of $\Psi(\sigma_{12})$ shows that the synchronized state is also unstable for the inter-layer link connecting oscillator number 2 from each layer. Finally, the green curve in Fig. 5 represents the case in which both oscillators 1 and 2 from each layer receive an inter-layer link. In this case, $\Psi(\sigma_{12})$ becomes negative around $\sigma_{12}=0.3$, indicating the onset of stable synchronized states. Therefore, stable complete synchronization among the oscillators from different layers is possible only when the number of inter-layer links equals the number of oscillators per layer.
@@
Thus, we confirm that a two-layer system with two oscillators per layer cannot achieve complete synchronization when connected by only a single inter-layer link. Next, we would like to consider dynamically switching this link between the two oscillators in each layer. However, it is important to note that such link switching may eventually introduce discontinuities in the system, raising questions about the well-definedness and convergence of $\Psi(\sigma_{12},T_{swt})$. To address this difficulty, we couple the oscillator via a smooth function that approximates a piecewise constant one, periodically alternating between $1$ and 00 within prescribed time intervals. This function $g\left(t,\alpha,p\right)$ is defined as:
@@
where the parameter $\alpha$ controls the sharpness of the transition between 00 and $1$, and $p$ determines the periodicity of $g(t,\alpha,p)$. The analysis of the synchronization errors shows that the results are coinciding for $\alpha\geq 5$. Therefore, $\alpha=5$ is used for all simulations. In this context, the period $p$ is twice the switching time, $T_{swt}$. The variable $n=\left\lfloor\frac{t}{p}\right\rfloor$ denotes the number of complete periods of length $p$ that fit within a given time interval. Consequently, $g\left(t,\alpha,p\right)$ generates a periodic pattern with intervals of length $p$, creating a smooth, square-wave-like shape. Therefore, this function offers a continuous method for switching an inter-layer link on and off, connecting a pair of mirror oscillators in a two-layer system containing two oscillators per layer. To simulate the switching of the inter-layer link between pairs of mirror oscillators, we connect one pair of mirror oscillators using a function $f\left(t,\alpha,p\right)=g(t+p,\alpha,p)$,
which is in the opposite phase of $g\left(t,\alpha,p\right)$ connecting the other pair. The schematic shown in Fig. 6(a) illustrates the connection between mirror oscillators number 2 by the function $g\left(t,\alpha,p\right)$ in blue, while the connection between mirror oscillator number 1 is mediated by the function $f\left(t,\alpha,p\right)$ in red. In addition, in Fig. 6(b), we illustrate the behavior of $g\left(t,\alpha,p\right)$ and $f\left(t,\alpha,p\right)$, simulating the switching of the inter-layer link between the two mirror oscillators.
@@
Figure 6: (a) Schematic of the two-layer system with two oscillators per layer, represented by blue and red circles. The intra-layer links are depicted in black. The inter-layer link connecting mirrored oscillators 1 is highlighted in red and mediated by the smooth square-wave function $f(t)$, while the inter-layer link connecting mirrored oscillators 2 is shown in blue and mediated by the smooth square-wave function $g(t)$. (b) The time evolution of the function $f(t)$ (red curve) and $g(t)$ (blue curve) is shown. The opposite phases between $f(t)$ and $g(t)$ illustrate the on/off dynamics of the inter-layer links. The parameters used to illustrate smooth switching are $\alpha=5$ and $p=50$, i.e. $T_{swt}=25$.
Because only one inter‑layer link is active at any instant, the inter‑layer Laplacian
$\mathcal{L}^{I}=\begin{pmatrix}-1&1\\[2.0pt] 1&-1\end{pmatrix}$
acts on a single pair of mirror nodes, while the coupling matrix $\Gamma(t)=\operatorname{diag}\bigl{(}\gamma_{1}(t),0,\dots,\gamma_{2N}(t),0\bigr{)}$ is diagonal. Hence, the intra‑ and inter‑layer operators commute,
$\bigl{[}\mathcal{L},\;\mathcal{L}^{I}\!\otimes\!\Gamma(t)\bigr{]}=0,$
so they are simultaneously diagonalizable (see [37],[38]). Diagonalising $\mathcal{L}^{I}$ yields two eigenvalues, $\lambda_{1}=0$ (longitudinal) and $\lambda_{2}=-2$ (transverse). Because $\Gamma(t)$ is diagonal, the spectrum of the Kronecker product factorises,
$\Lambda\!\bigl{(}\mathcal{L}^{I}\!\otimes\!\Gamma(t)\bigr{)}=\{0,\,-2\}\times\{\gamma_{j}(t)\}_{j=1}^{2N},$
so the set of eigenvalues is independent of the layer size $N$; $N$ affects only the multiplicities. Consequently, the full $2N$‑node duplex has a single non‑trivial transverse block of dimension $2$,
[…]
whose largest Lyapunov exponent is the master‑stability function $\Psi(\sigma_{12},T_{\text{swt}})$. If this block is stable, every duplicate block is equally stable, and the synchronization manifold is therefore stable for all $N$.
Limitations: The reduction relies on
  1. i
    a single switched inter‑layer link and
  2. ii
    a diagonal $\Gamma(t)$. Switching several links simultaneously or coupling both $u$ and $v$ variables would violate the commutation property and require a higher‑dimensional stability analysis.
@@
Next, we perform a transverse stability analysis using the MSF for the system with the continuous switching of the inter-layer link. Hence, in Fig. 7, we present $\Psi(\sigma_{12})$ as a function of coupling intensity $\sigma_{12}$ for various switching times $T_{swt}$ (indicated by different colored curves). For $T_{swt}=135$ (purple curve in Fig. 7), we observe that $\Psi(\sigma_{12})$ remains positive across the entire range of $\sigma_{12}$, indicating that stable, fully synchronized states do not occur, even at higher values of $\sigma_{12}$. However, at a shorter switching time, $T_{swt}=100$, we find that $\Psi(\sigma_{12})$ crosses the zero axis at approximately $\sigma_{12}=0.9$ (red curve in Fig. 7), signaling the onset of stable, fully synchronized states. As we further decrease the values of $T_{swt}$, the onset of fully synchronized states occurs at progressively smaller values of $\sigma_{12}$. This is illustrated in Fig. 7, where the intercepts of the green and yellow curves correspond to $T_{swt}=50$ and $T_{swt}=25$, respectively. Finally, for $T_{swt}=11$, the corresponding curve for $\Psi(\sigma_{12})$ crosses the zero axis at the smallest observed value of $\sigma_{12}$. Thus, decreasing $T_{swt}$ further does not facilitate the onset of fully synchronized states at even lower values of $\sigma_{12}$.
@@
Furthermore, the results in Figs. 8(a) and 8(b) effectively illustrate the relationship between $T_{swt}$ and $\sigma_{12}$. Specifically, they clarify that below the switching time threshold of approximately $T_{swt}=40$, reducing $T_{swt}$ does not enable synchronization at a coupling intensity lower than about $\sigma_{12}\approx 0.33$.

### v2

@@ III.1 Link-switching-induced inter-layer synchronization
We now demonstrate the role of switching the existing inter-layer links in establishing synchronization between the layers. Naturally, the time interval during which the links remain active before switching, referred to as the switching time $T_{swt}$, plays a central role in this process. To explore this role, we first analyze the time evolution of the double-layer system for different values of $T_{swt}$ in a network with $N=400$ FHN oscillators per layer and only $N_{IL}=100$ inter-layer links connecting the mirror nodes between the layers.The parameters used for the figures are all common and as following: $\varepsilon=0.05$, $a_{ij}=0.87$ (even $j$), $a_{ij}=0.97$ (odd $j$), $\phi=\frac{\pi}{2}-0.1$, $\sigma_{1}=\sigma_{2}=0.1$.In Fig. 3(a), for $T_{swt}=120$, the red curve represents the time evolution of the Euclidean distance between mirror nodes numbered $35$, defined as $E_{35}=\sqrt{(u_{i=1j=35}-u_{i=2j=35})^{2}+(v_{i=1j=35}-v_{i=2j=35})^{2}}$. In this figure, the time intervals shown in gray correspond to the presence of an inter-layer link between the mirror nodes $35$, while the time intervals shown in white indicate the absence of such a link. We observe that the distance $E_{35}$ exhibits high-amplitude oscillations during the time intervals marked in white, corresponding to the absence of an inter-layer link. In contrast, during the gray intervals, where an inter-layer link is present, $E_{35}$ oscillates with a significantly lower amplitude. Notice that, for $T_{swt}=120$, immediately after the gray time intervals, regardless of their duration, the amplitude of $E_{35}$ begins to increase as soon as the inter-layer link between mirror nodes $35$ is switched off. This behavior suggests that switching the inter-layer links at every $T_{swt}=120$ does not result in inter-layer synchronization. On the other hand, in Fig. 3(b), for $T_{swt}=23$, where the resulting time intervals marked in gray are significantly shorter, we observe a transient phase in which high and low amplitude oscillations alternate depending on the presence or absence of an inter-layer link. However, in this case, right after one gray time interval, the distance $E_{35}$ drops to zero, marking the end of the transient phase and indicating synchronization between the mirror nodes $35$. Finally, we emphasize that, although all links are switched periodically at every $T_{swt}$, a link connecting a specific pair of mirror nodes may be reassigned to the same pair after switching. Consequently, as observed in Fig. 3(a) and Fig. 3(b), the time intervals marked in gray can vary in size. Moreover, since $N_{IL}=0.25N$, a given pair of mirror nodes may not receive an inter-layer link even after several switches, resulting in the time intervals marked in white also varying in size.Now, we complement the analysis presented in Figs. 3(a) and 3(b) by verifying their corresponding behavior with respect to inter-layer synchronization. To achieve this, we first define a measure to estimate the level of inter-layer synchronization at each time instant as:
$E^{12}(t)=\frac{1}{N}\|\mathbf{L}_{1}(t)-\mathbf{L}_{2}(t)\|,$ (4)
@@
Figure 3 illustrates the dynamics of the double-layer system influenced by the switching of inter-layer links. We observe that switching the existing links at every $T_{swt}=23$ drives the system to inter-layer synchronization. To verify the robustness of this switching process, we evaluate the duration of the transient time required for the link switching to bring the layers into complete inter-layer synchronization.In Fig. 4, we present the duration of the transient time required for the link switching to establish complete inter-layer synchronization, $T_{sync}$, as a function of the layer size $N$ for different switching times $T_{swt}=5,10,15,23,25,35,50$, while keeping the ratio $N_{IL}/N=0.25$ constant complying with the sparse network requirement.The transient time $T_{sync}$ increases monotonically with increasing $N$ for all switching times, demonstrating the effectiveness of the link switching procedure.
@@
Concerning the dependence of effectiveness of this link-switching strategy on the system parameters, we expect the qualitative effectiveness to persist across different chaotic attractors. Nevertheless, chaotic attractors may differ substantially in the structure of their tangent spaces, including the magnitude and distribution of local contraction and expansion rates. As a result, the quantitative effectiveness of the link-switching strategy may depend on specific features of the underlying chaotic dynamics.
@@ III.2 Stability analysis of synchronized states
In the previous section, we observed that with a low density of inter-layer links, the dynamic redistribution of existing links among the FHN oscillators can promote the emergence of synchronous dynamics across the layers. More interestingly, the timing of these redistributions, defined by the switching time $T_{swt}$, plays a crucial role. However, whether the emergent synchronous dynamics is asymptotically stable remains an open question. To address this issue, we perform a transverse stability analysis using the master stability function (MSF) formalism on a simplified system consisting of two FHN oscillators per layer.To implement the MSF in such a system, we treat each identical layer, comprising two oscillators, as a single higher-dimensional dynamical system, rather than as a network of lower-dimensional systems. Accordingly, we define the function $\mathbf{G}(\mathbf{L}_{i}):=\mathbf{F}(\mathbf{L}_{i})+\sigma_{i}\,(\mathcal{L}\!\otimes\!\mathcal{H})\,\mathbf{L}_{i}$, with $i=1,2$ and $\sigma\equiv\sigma_{1}=\sigma_{2}$, which encapsulates the local dynamics of the FHN oscillators defined in Eq. (2) and the intra-layer coupling prescribed by the matrices $\mathcal{L}$ and $\mathcal{H}$, together with the parameter $\sigma_{i}$. Using this approach, Eq. (1) can be rewritten as:
@@
where $\mathbf{L}_{i}$ continue to represent the dynamical state of the entire layer, which now comprises $N=2$ FHN oscillators labeled as $j=1,2$.The complete synchronized inter-layer state in the double-layer system occurs on a synchronization manifold $\mathbf{S}$ which is defined as $\mathbf{S}=\mathbf{L}_{1}=\mathbf{L}_{2}$. To assess the local stability of $\mathbf{S}$ using the MSF formalism, we introduce infinitesimal perturbations $\delta\mathbf{L}_{i}$ in all directions of the layer state space. This approach leads to the following variational equation governing the time evolution of $\delta\mathbf{L}_{i}$:
$\dot{\delta{\bf L\;}}_{i}=\left[\mathbf{DG}(\mathbf{S})+\sigma_{12}(\mathcal{L}^{I}\otimes\mathbf{\Gamma}(t))\right]\cdot\delta\mathbf{L}_{i},$ (6)
where $\mathbf{DG}(\mathbf{S})$ represent the Jacobian matrices of the intra-layer dynamics evaluated at the synchronization manifold $\mathbf{S}$. The Laplacian matrix $\mathcal{L}^{I}$ is diagonalized with real eigenvalues $\lambda_{i}$ for $i=1,2$. Consequently, Eq. (6) becomes block diagonalizable, yielding a decoupled variational equation:
$\delta\dot{\mathbf{\Theta}}_{i}=\left[\mathbf{DG}(\mathbf{S})-\sigma_{12}|\lambda_{i}|\mathbf{\Gamma}(t)\right]\cdot\delta\mathbf{\Theta}_{i},$ (7)
@@
$\delta\dot{\mathbf{\Theta}}=\left[\mathbf{DG}(\mathbf{S})-2\sigma_{12}\mathbf{\Gamma}(t)\right]\cdot\delta\mathbf{\Theta}.$ (8)
The largest Lyapunov exponent of the system in Eq. (8) can be represented by the MSF $\Psi$. Since the coupling intensity $\sigma_{12}$ and the switching time $T_{swt}$ are the parameters of interest, we obtain the MSF as function of these parameters $\Psi(\sigma_{12},T_{swt})$.Now, to demonstrate that a single inter-layer link is insufficient to establish synchronization between the oscillators of the two-layer system, we calculate $\Psi(\sigma_{12},T_{swt})$ as a function of $\sigma_{12}$ for the static case, i.e., without link switching. Hence, in Fig. 6, the blue curve illustrates the behavior of $\Psi(\sigma_{12})$ for the inter-layer link connecting oscillator number 1 from each layer (see schematic in Fig. 7(a) for the oscillators numbers). In this case, $\Psi(\sigma_{12})$ is positive within the considered interval of $\sigma_{12}$, indicating that the completely synchronized state is unstable for this inter-layer topology. Similarly, the red curve on the positive side of $\Psi(\sigma_{12})$ shows that the synchronized state is also unstable for the inter-layer link connecting oscillator number 2 from each layer. Finally, the green curve in Fig. 6 represents the case in which both oscillators 1 and 2 from each layer receive an inter-layer link. In this case, $\Psi(\sigma_{12})$ becomes negative around $\sigma_{12}=0.15$, indicating the onset of stable synchronized states. Therefore, stable complete synchronization among the oscillators from different layers is possible only when the number of inter-layer links equals the number of oscillators per layer.
@@
Thus, we confirm that a two-layer system with two oscillators per layer cannot achieve complete synchronization when connected by only a single inter-layer link. Next, we would like to consider dynamically switching this link between the two oscillators in each layer. However, it is important to note that such link switching may eventually introduce discontinuities in the system, raising questions about the well-definedness and convergence of $\Psi(\sigma_{12},T_{swt})$.To address this difficulty, we couple the oscillator via a smooth function that approximates a piecewise constant one, periodically alternating between $1$ and 0 within prescribed time intervals. This function $g\left(t,\alpha,p\right)$ is defined as:
@@
where the parameter $\alpha$ controls the sharpness of the transition between 0 and $1$, and $p$ determines the periodicity of $g(t,\alpha,p)$. The analysis of the synchronization errors shows that the results are coinciding for $\alpha\geq 5$. Therefore, $\alpha=5$ is used for all simulations. In this context, the period $p$ is twice the switching time, $T_{swt}$. The variable $n=\left\lfloor\frac{t}{p}\right\rfloor$ denotes the number of complete periods of length $p$ that fit within a given time interval. Consequently, $g\left(t,\alpha,p\right)$ generates a periodic pattern with intervals of length $p$, creating a smooth, square-wave-like shape. Therefore, this function offers a continuous method for switching an inter-layer link on and off, connecting a pair of mirror oscillators in a two-layer system containing two oscillators per layer. To simulate the switching of the inter-layer link between pairs of mirror oscillators, we connect one pair of mirror oscillators using a function $f\left(t,\alpha,p\right)=g(t+p,\alpha,p)$, which is in the opposite phase of $g\left(t,\alpha,p\right)$ connecting the other pair. The schematic shown in Fig. 7(a) illustrates the connection between mirror oscillators number 2 by the function $g\left(t,\alpha,p\right)$ in blue, while the connection between mirror oscillator number 1 is mediated by the function $f\left(t,\alpha,p\right)$ in red. In addition, in Fig. 7(b), we illustrate the behavior of $g\left(t,\alpha,p\right)$ and $f\left(t,\alpha,p\right)$, simulating the switching of the inter-layer link between the two mirror oscillators.
@@
Figure 7: (a) Schematic of the two-layer system with two oscillators per layer, represented by blue and red circles. The intra-layer links are depicted in black. The inter-layer link connecting mirrored oscillators 1 is highlighted in red and mediated by the smooth square-wave function $f(t)$, while the inter-layer link connecting mirrored oscillators 2 is shown in blue and mediated by the smooth square-wave function $g(t)$. (b) The time evolution of the function $f(t)$ (red curve) and $g(t)$ (blue curve) is shown. The opposite phases between $f(t)$ and $g(t)$ illustrate the on/off dynamics of the inter-layer links.The parameters used to illustrate smooth switching are $\alpha=5$ and $p=25$, i.e. $T_{swt}=12.5$.
@@
Next, we perform a transverse stability analysis using the MSF for the system with such continuous switching of the inter-layer link. Hence, in Fig. 8, we present $\Psi(\sigma_{12})$ as a function of coupling intensity $\sigma_{12}$ for various switching times $T_{swt}$ (indicated by different colored curves). For $T_{swt}=135$ (purple curve in Fig. 8), we observe that $\Psi(\sigma_{12})$ remains positive across the entire range of $\sigma_{12}$, indicating that stable, fully synchronized states do not occur, even at higher values of $\sigma_{12}$. However, at a shorter switching time, $T_{swt}=100$, we find that $\Psi(\sigma_{12})$ crosses the zero axis at approximately $\sigma_{12}=0.9$ (red curve in Fig. 8), signaling the onset of stable, fully synchronized states. As we further decrease the values of $T_{swt}$, the onset of fully synchronized states occurs at progressively smaller values of $\sigma_{12}$. This is illustrated in Fig. 8, where the intercepts of the green and yellow curves correspond to $T_{swt}=50$ and $T_{swt}=25$, respectively. Finally, for $T_{swt}=11$, the corresponding curve for $\Psi(\sigma_{12})$ crosses the zero axis at the smallest observed value of $\sigma_{12}$. Thus, decreasing $T_{swt}$ further does not facilitate the onset of fully synchronized states at even lower values of $\sigma_{12}$. In addition, by comparing the stability results in Fig. 8 with those in Fig. 6 for the static inter-layer topology, we point out that fast switching in Fig. 8 effectively mimics the stability case in which both inter-layer links are connected in the static topology (green curve in Fig. 6), whereas slow switching tends toward the unstable case with only a single static inter-layer link (red and blue curves in Fig. 6).
@@
Furthermore, the results in Figs. 9(a) and 9(b) effectively illustrate the relationship between $T_{swt}$ and $\sigma_{12}$. Specifically, they clarify that below the switching time threshold of approximately $T_{swt}=40$, reducing $T_{swt}$ does not enable synchronization at a coupling intensity lower than about $\sigma_{12}\approx 0.33$.